The length of a rectangular football field is 5 yards less than twice the width. If the perimeter of the field is 380 yards, find the dimensions of the football field.

Answer by checkley79(3341)

P=2L+2W
L=2W-5
380=2(2W-5)+2W
380=4W-10+2W
380=6W-10
6W=380+10
6W=390
W=390/6
W=65 ANS. FOR THE WIDTH.
L=2*65-5
L=130-5
L=125 ANS. FOR THE LENGTH.
PROOF:
380=2*125+2*65
380=250+130
380=380